jennylove:

If 20 waves of water pass you in 5.0 s and the waves are 2.0 m long, how fast are they moving? 1.0 m/s 2.0 m/s 4.0 m/s 8.0 m/s

jennylove:

2.0 m/s ?

justjm:

\(v=λƒ\) λ = 2.0 m ƒ = 20 waves/5.0 seconds = 4.0 hz hence 8 I believe

jennylove:

how did you know that it would be 8 with the 4.0 hz?

jennylove:

if i helps, this one was C (4.0) If 20 waves of water pass you in 10.0 s and the waves are 2.0 m long, how fast are they moving? 1.0 m/s 2.0 m/s 4.0 m/s 8.0 m/s

justjm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jennylove how did you know that it would be 8 with the 4.0 hz? \(\color{#0cbb34}{\text{End of Quote}}\) It is asking for velocity of the wave, so you must take the frequency, 4.0 hz, then multiply that with wavelength, 2 m

justjm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jennylove if i helps, this one was C (4.0) If 20 waves of water pass you in 10.0 s and the waves are 2.0 m long, how fast are they moving? 1.0 m/s 2.0 m/s 4.0 m/s 8.0 m/s \(\color{#0cbb34}{\text{End of Quote}}\) Yes it is 4. Frequency is 20/10 = 2 hz, and wavelength = 2.0, so v=2*2=4

jennylove:

Oh okay, that makes sense. I was way off lol, thanks

justjm:

I think you got 2.0 m/s using the other formula for velocity of a wave, v=\(\frac{λ}{T}\)